OpenStudy (pinkkjazzy3):

solve for w. w/4=w/5-3 simplify.

jabez177 (jabez177):

\[\frac{ w }{ 4 }=\frac{ w }{ 5-3 }\]Right? :/

OpenStudy (pinkkjazzy3):

yes

jhonyy9 (jhonyy9):

i think 5-3 like denominator ot make sense hope may be right w/4 = (w/5) - 3 and hence will get like common denominator the 20 and will result 5w = 4w - 60 5w-4w = -60 w = -60 hope helped